[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed May 24, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Wed May 24 17:45:56 UTC 2017


Here is today's summary of NHC computer operations:
--- Mello

----------

TAFB:
Andy Latto's low resolution web images are updating again.


TAFB Color Laser Jet Printer (large printer in TAFB ops) still needs to be  
moved to the copy room. The front office along with Hugh and Mark  
coordinated this yesterday. A NCO ticket remains open for the task.

Awips2 bad NAS disks replaced.
HSU:
None
Other items:

NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170524/0f83fc00/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list